# SPECIAL RELATIVITY

### v9

@@ 13.4 The Usual Integration of the 4D Covariant Equation of Motion
We should then expect to get results similar to those obtained in the case of the (3+1) non-covariant particle tracking. In fact, based on the structure of the four components of the equation of motion Eq.(20), we can arrive to another mathematically identical formulation of the dynamical problem. The fact that the evolution of the particle in the lab frame is subject to a constraint has already been mentioned. This means that the mathematical form of the dynamics law includes only three independent equations of motion. It is easy to see from the initial set of four equations, $du^{(0)}/d\tau=0$, $du^{(1)}/d\tau=0$, $du^{(2)}/d\tau=-eBu^{(3)}/(mc)$, $du^{(3)}/d\tau=eBu^{(2)}/(mc)$, that the presentation of the time component simply as the relation $d\tau=dt/\gamma$ between proper time and coordinate time is just a simple parametrization that yields the corrected Newton’s equation Eq.(21) as another equivalent form of these four equations in terms of absolute time $t$ instead of proper time of the particle. This approach to integrating dynamics equations from the initial conditions relies on the use of three independent spatial coordinates and velocities without constraint and is intimately connected with old kinematics. The presentation of the time component simply as the relation $d\tau=dt/\gamma$ between proper time and coordinate time is based on the hidden assumption that the type of clock synchronization, which provides the time coordinate $t$ in the lab frame, is based on the use of the absolute time convention.
@@ 13.5 Covariant Particle Tracking
In conventional particle tracking a particle trajectory $\vec{x}(t)$ can be seen from the lab frame as the result of successive Galileo boosts that track the motion of the accelerated (in a constant magnetic field) particle. The usual Galileo rule for addition of velocities is used to determine the Galileo boosts tracking a particular particle, instant after instant, along its motion along the curved trajectory.
@@
We cannot take old Newton’s kinematics for mechanics and Einstein’s kinematics for electrodynamics. Only solution of the dynamics equations in covariant form gives the correct coupling between the usual Maxwell’s equations and particle traectories in the lab frame. The trajectory $\vec{x}_{cov}(t)$ is viewed from the lab frame as the result of successive Lorentz transformations. In this case relativistic kinematics effects arise. Let us try out our algorithm for reconstructing $\vec{x}_{cov}(t)$ on some example, to see how it works. We will find formulas relating the velocity of a particle in one reference system to its velocity in a second reference system. First, we examine the reasoning presented in textbooks. Let us suppose that the $K^{\prime}$ system moves relative to the $K$ system with velocity $V$ along the $x$ axis. Let $\vec{v}=d\vec{x}/dt,$ be the vector of the particle velocity in the $K$ system and $\vec{v^{\prime}}=d\vec{x^{\prime}}/dt^{\prime}$ the velocity vector of the same particle in the $K^{\prime}$ system. From Lorentz transformation we have $dx=\gamma(dx^{\prime}+Vdt^{\prime})$, $dy=dy^{\prime}$, $dz=dz^{\prime}$, $dt=\gamma(dt^{\prime}+Vdx^{\prime}/c^{2})$, where $\gamma=1/\sqrt{1-V^{2}/c^{2}}$. Dividing the first three equations by the fourth we find $v_{x}=(v_{x}^{\prime}+V)/(1+Vv^{\prime}_{x}/c^{2})$, $v_{y}=v^{\prime}_{y}/[\gamma(1+Vv^{\prime}_{x}/c^{2})]$, $v_{z}=v^{\prime}_{z}/[\gamma(1+Vv^{\prime}_{x}/c^{2})]$. These formulae determine the transformation of velocities. According to textbooks, they describe the law of composition of velocities in the theory of relativity.
@@
We begin with the simple case of relativistically combining perpendicular velocities. There is no length contraction for perpendicular distances, but time dilation still occur. Therefore the particle velocity in the frame $K$ is just $v_{x}=V$, $v_{y}=v_{y}^{\prime}/\gamma$. Note that we must be clear what we mean by ”perpendicular”. It only makes sense to say two velocities are perpendicular if they are measured in the same reference frame. Hence when authors of textbooks say that $\vec{e}_{x}^{\prime}v_{x}^{\prime}$ and $\vec{e}_{x}V$ are perpendicular they actually mean that $(x^{\prime},y^{\prime},z^{\prime})$ axes and $(x,y,z)$ axes are parallel. In other words, it is (hidden) assumed that $K$ and $K^{\prime}$ observers have common 3-space. This is misconception.
@@
One of the consequences of non-commutativity of non-collinear Lorentz boosts is not the existence of a common ordinary space. Suppose that in the Lorentz frame $K^{\prime}$ the traveler was observing particle motion along the $y^{\prime}$ axis. In other words measured velocity of the particle $v^{\prime}_{y}$ , and yet the frame is moving relative to the frame $K$ along $x$ axis. Then to an observer at rest in the Lorentz frame $K$, the frame $K^{\prime}$ with the moving particle seems to rotate about the frame $K$. Let us consider the simple case when the velocity $V$ approaches to that of light and the velocity $v^{\prime}_{y}$ is small. This means that we take limit $\gamma\gg 1\gg v^{\prime}_{y}/c$. We consider the small parameter $v^{\prime}_{y}/c$ and use the second order approximation. For ultrarelativistic limit $\gamma\gg 1$ and the axes of $K^{\prime}$ frame are then rotated with respect to the $K$ frame axes by angle $\theta_{w}=v^{\prime}_{y}/(\gamma c)$. We want to look now at the consequences of this rotation. First we notice that there is a projection of the velocity $v^{\prime}_{y}$ on the $x$-axis, so the velocity component along the $x$-axis in the frame $K$ is going to be smaller than $V$. Our problem now is to work out the decrement of the horizontal velocity in the $K$ frame. How we can do that? First, the relativistic correction to the horizontal velocity appears only in the order $[v^{\prime}_{y}/(\gamma c)]^{2}$. The trajectory of the particle is viewed from the $K$ frame as the result of successive infinitesimal Lorentz transformations. Integrating with respect to infinitesimal value $dv^{\prime}_{y}$, we find that the horizontal velocity is $v_{x}=V-(v^{\prime}_{y})^{2}/(2\gamma^{2}c)$. In the ultrarelativistic approximation we find the simple result $v_{y}=v^{\prime}_{y}/\gamma$, $v_{x}=V-v_{y}^{2}/(2c)$, so that Lorentz transformations simply leads to a rotation of particle velocity $V$ of the angle $v_{y}/c=v_{y}/V$. In contrast, according to textbooks, it follows that the total particle speed in the Lorentz $K$ frame increase from $V$ to $V+v_{y}^{2}/(2c)$. Our result is at odds with the prediction from textbooks. The commonly accepted derivation of the composition of velocities based on the use the relation $dt^{\prime}=dt/\gamma$ and does not account for the mixture of (transverse) positions and time.
Now we go on to consider another situation, a very practical one. We want to study the case of relativistically combining perpendicular velocities where velocity $v_{y}^{\prime}$ approaches to that of light, $v_{y}^{\prime}\to c$, and the velocity $V$ is small, $V/c\ll 1$. According to textbooks, the velocity addition formula gives $v_{x}\to V$, $v_{y}=v_{y}^{\prime}/\gamma\to v_{y}^{\prime}-V^{2}/(2c)$. So that two Lorentz boosts simply leads to a rotation of particle velocity $v_{y}^{\prime}$ of the angle $V/c$. We can believed that it is right; it is example what called aberration of (ultra) relativistic particles and it is well known. How can that be? Perhaps it is not puzzling, through, when we remember that what we called inertial frame $K^{\prime}$ is really accelerated frame with respect to the inertial lab frame $K$. It is important to stress in this point that we consider the motion of a particle in the lab frame and we want to describe how to determine the covariant trajectory $\vec{x}(t)$. Above in the Chapter 7 we demonstrated that the aberration of particles in the accelerated frame $K^{\prime}$ can be described in the lab frame within standard special relativity taking advantage of the Wigner rotation theory. As viewed from the lab frame, the coordinate axes of the accelerated frame is rotated through the angle $V/c$ with respect to the coordinate axes of the lab frame. We can easily understand that the particle motion in the $K^{\prime}$ frame will be parallel to the $y$ coordinate axis of the lab frame. It is clear that we should then expect to get results similar to those obtained in the textbooks. When we studied asymptotic $V\to c$, $v_{y}^{\prime}/c\ll 1$, we saw that when particle is moving nonrelativistically in the frame $K^{\prime}$ (i.e. when the aberration effect is negligible), a particle motion will be parallel to the coordinate axis $y^{\prime}$ and we have no problem of demonstrating that there is the connection between the velocity transformation and the Wigner rotation.
It is hoped that presented example will draw wide attention to the central role that the Wigner rotation plays in the transformation of non-collinear velocities. To emphasize the physical concepts involved, rather than mathematical formalism, we consider not the most general case of the (corrected) velocity addition theorem. However, the special case considered already involves all necessary ingredients.
@@ 13.6 On the Relativistic Velocity Composition and the Wigner Rotation
There are two difficulties in solving problems concerning covariant particle tracking. First of all, conventional particle tracking in the single inertial frame is actually based on classical Newton mechanics without reference to Lorentz transformations. According to textbooks, if no more than one frame is involved, one does not need to know the theory of relativity. In other words, in this standard approach, introducing as only modification to the classical case the relativistic mass, time differ from space. It fact, we have no mixture positions and time. This essential point has never received attention in the physical community.
Another difficulty is the confusion surrounding the combination of non-parallel velocities in special relativity. Let us to consider composition of two velocities when they are not lying on a straight line. In conventional particle tracking (similar to the Newton mechanics), one composes them geometrically by the method of parallelogram with velocities regarded as the usual vectors. In relativistic kinematics, the composition of the velocities is studied analytically from the composition of Lorentz boosts. At the same time it is common knowledge that Lorentz boosts alone do not make a group. We must conclude that when we are dealing with transformations and set of these transformations does not form a group we have no legal theory. Authors of textbooks believe that the relativity principle requires that Einstein’s velocity addition law gives the composition law of a group. The non-associativity of the composition law of the non-colinear velocities is not widely known. Many experts who learned the theory of relativity using textbooks will find this statement disturbing. The popular belief remains that Einstein’s law of addition of velocities is associative, although it is an exercise in elementary algebra to prove non-associativity (3). It is therefore impossible to get covariant trajectory, $\vec{x}_{cov}(t)$, if we have nothing but Lorentz boosts. Something else has to be added. Lorentz boosts do not form a group due to presence of the Wigner rotation. It is comes out quite natural that breakdown of commutativity and associativity in Einstein addition, turn out to be Wigner rotation.
@@ 13.7 Convention-Invariant Particle Tracking
So far we have considered the motion of a particle in three-dimensional space using the vector-valued function $\vec{x}(t)$. We have a prescribed curve (path) along which the particle moves. The motion along the path is described by $l(t)$, where $l$ is a certain parameter (in our case of interest the length of the arc). Note the difference between the notions of path and trajectory. The trajectory of a particle conveys more information about its motion because every position is described additionally by the corresponding time instant. The path is rather a purely geometrical notion. Complete paths or their parts may consist of, e. g., line segments, arcs, circles, helical curves. If we take the origin of the (Cartesian) coordinate system and we connect the point to the point laying on the path and describing the motion of the particle, then the creating vector will be a position vector $\vec{x}(l)$. The derivative of a vector is the vector tangent to the curve described by the radius vector $\vec{x}(l)$. The sense of the $d\vec{x}(l)/dl$ is determined by the sense of the curve arc $l$.
We already know from our discussion in Introduction that the path $\vec{x}(l)$ has exact objective meaning i.e. it is convention-invariant. The components of the momentum four vector $mu=(\mathcal{E}/c,\vec{p})$ have also exact objective meaning. In contrast to this, and consistently with the conventionality intrinsic in the velocity, the trajectory $\vec{x}(t)$ of the particle in the lab frame is convention dependent and has no exact objective meaning.
@@
These three equations corresponds exactly to the equations for the components of the position vector that can be found using the non-covariant particle tracking approach, and $\vec{x}(l)_{cov}$ is exactly equal to $\vec{x}(l)$ as it must be. The point is that both approaches describe correctly the same physical reality and since the curvature radius of the path in the magnetic field, and consequently the three-momentum, has obviously an objective meaning (i.e. is convention-invariant), both approaches yield the same physical results.
13.8 The Nature of Mass
This section concentrates on the nature of mass. The concept of mass is one of the most fundamental notions in physics, comparable in importance only to concept of space and time. With advent of the special theory of relativity, physicists and philosophers focused their attention on the concept of relativistic mass. The remarkable progress in experimental and theoretical physics made during the past few decades has considerably deepened our knowledge concerning the nature of mass. We conclude this section with a brief discussion of the physical meaning of the Minkowski geometry.
13.8.1 Phenomenology and Relativistic Extensions

### v10

@@ 16.4 The Usual Integration of the 4D Covariant Equation of Motion
We should then expect to get results similar to those obtained in the case of the (3+1) non-covariant particle tracking. In fact, based on the structure of the four components of the equation of motion Eq.(21), we can arrive at another mathematically identical formulation of the dynamical problem. The fact that the evolution of the particle in the lab frame is subject to a constraint has already been mentioned. This means that the mathematical form of the dynamics law includes only three independent equations of motion. It is easy to see from the initial set of four equations, $du^{(0)}/d\tau=0$, $du^{(1)}/d\tau=0$, $du^{(2)}/d\tau=-eBu^{(3)}/(mc)$, $du^{(3)}/d\tau=eBu^{(2)}/(mc)$, that the presentation of the time component simply as the relation $d\tau=dt/\gamma$ between proper time and coordinate time is just a simple parametrization that yields the corrected Newton’s equation Eq.(22) as another equivalent form of these four equations in terms of absolute time $t$ instead of proper time of the particle. This approach to integrating dynamics equations from the initial conditions relies on the use of three independent spatial coordinates and velocities without constraint and is intimately connected with old kinematics. The presentation of the time component simply as the relation $d\tau=dt/\gamma$ between proper time and coordinate time is based on the hidden assumption that the type of clock synchronization, which provides the time coordinate $t$ in the lab frame, is based on the use of the absolute time convention.
@@ 16.5 Covariant Particle Tracking
In conventional particle tracking a particle trajectory $\vec{x}(t)$ can be seen from the lab frame as the result of successive Galileo boosts that track the motion of the accelerated (in a constant magnetic field) particle. The usual Galileo rule for the addition of velocities is used to determine the Galileo boosts tracking a particular particle, instant after instant, along its motion along the curved trajectory.
@@
We cannot take old Newton’s kinematics for mechanics and Einstein’s kinematics for electrodynamics. Only the solution of the dynamics equations in covariant form gives the correct coupling between the Maxwell’s equations and particle trajectories in the lab frame. The trajectory $\vec{x}_{cov}(t)$ is viewed from the lab frame as the result of successive Lorentz transformations. In this case, relativistic kinematics effects arise. Let us try out our algorithm for reconstructing $\vec{x}_{cov}(t)$ on some examples, to see how it works. We will find formulas relating the velocity of a particle in one reference system to its velocity in a second reference system. First, we examine the reasoning presented in textbooks. Let us suppose that the $K^{\prime}$ system moves relative to the $K$ system with velocity $V$ along the $x$ axis. Let $\vec{v}=d\vec{x}/dt,$ be the vector of the particle velocity in the $K$ system and $\vec{v^{\prime}}=d\vec{x^{\prime}}/dt^{\prime}$ the velocity vector of the same particle in the $K^{\prime}$ system. From Lorentz transformation we have $dx=\gamma(dx^{\prime}+Vdt^{\prime})$, $dy=dy^{\prime}$, $dz=dz^{\prime}$, $dt=\gamma(dt^{\prime}+Vdx^{\prime}/c^{2})$, where $\gamma=1/\sqrt{1-V^{2}/c^{2}}$. Dividing the first three equations by the fourth we find $v_{x}=(v_{x}^{\prime}+V)/(1+Vv^{\prime}_{x}/c^{2})$, $v_{y}=v^{\prime}_{y}/[\gamma(1+Vv^{\prime}_{x}/c^{2})]$, $v_{z}=v^{\prime}_{z}/[\gamma(1+Vv^{\prime}_{x}/c^{2})]$. These formulae determine the transformation of velocities. According to textbooks, they describe the law of composition of velocities in the theory of relativity.
@@
We begin with the simple case of relativistically combining perpendicular velocities. There is no length contraction for perpendicular distances, but time dilation still occurs. Therefore the particle velocity in the frame $K$ is just $v_{x}=V$, $v_{y}=v_{y}^{\prime}/\gamma$. Note that we must be clear what we mean by ”perpendicular”. It only makes sense to say two velocities are perpendicular if they are measured in the same reference frame. Hence when authors of textbooks say that $\vec{e}_{y^{\prime}}v_{y}^{\prime}$ and $\vec{e}_{x}V$ are perpendicular they actually mean that $(x^{\prime},y^{\prime},z^{\prime})$ axes and $(x,y,z)$ axes are parallel. In other words, it is (hidden) assumed that $K$ and $K^{\prime}$ observers have common a 3-space. This is a misconception.
@@
One of the consequences of non-commutativity of non-collinear Lorentz boosts is not the existence of a common ordinary space. Suppose that in the Lorentz frame $K^{\prime}$ the traveler was observing particle motion along the $y^{\prime}$ axis. In other words, the measured velocity of the particle $v^{\prime}_{y}$ , and yet the frame is moving relative to the frame $K$ along $x$ axis. Then to an observer at rest in the Lorentz frame $K$, the frame $K^{\prime}$ with the moving particle seems to rotate about the frame $K$. Let us consider the simple case when the velocity $V$ approaches to that of light and the velocity $v^{\prime}_{y}$ is small. This means that we take limit $\gamma\gg 1\gg v^{\prime}_{y}/c$. We consider the small parameter $v^{\prime}_{y}/c$ and use the second-order approximation. For ultrarelativistic limit $\gamma\gg 1$ and the axes of $K^{\prime}$ frame are then rotated with respect to the $K$ frame axes by angle $\theta_{w}=v^{\prime}_{y}/(\gamma c)$. We want to look now at the consequences of this rotation. First, we notice that there is a projection of the velocity $v^{\prime}_{y}$ on the $x$-axis, so the velocity component along the $x$-axis in the frame $K$ is going to be smaller than $V$. Our problem now is to work out the decrement of the horizontal velocity in the $K$ frame. How we can do that? First, the relativistic correction to the horizontal velocity appears only in the order $[v^{\prime}_{y}/(\gamma c)]^{2}$. The trajectory of the particle is viewed from the $K$ frame as the result of successive infinitesimal Lorentz transformations. Integrating with respect to infinitesimal value $dv^{\prime}_{y}$, we find that the horizontal velocity is $v_{x}=V-(v^{\prime}_{y})^{2}/(2\gamma^{2}c)$. In the ultrarelativistic approximation we find the simple result $v_{y}=v^{\prime}_{y}/\gamma$, $v_{x}=V-v_{y}^{2}/(2c)$, so that Lorentz transformations simply lead to a rotation of particle velocity $V$ of the angle $v_{y}/c=v_{y}/V$. In contrast, according to textbooks, it follows that the total particle speed in the Lorentz $K$ frame increases from $V$ to $V+v_{y}^{2}/(2c)$. Our result is at odds with the prediction from textbooks. The commonly accepted derivation of the composition of velocities is based on the use of the relation $dt^{\prime}=dt/\gamma$ and does not account for the mixture of (transverse) positions and time.
@@ 16.6 On the Relativistic Velocity Composition and the Wigner Rotation
There are two difficulties in solving problems concerning covariant particle tracking. First of all, conventional particle tracking in the single inertial frame is based on classical Newton mechanics without reference to Lorentz transformations. According to textbooks, if no more than one frame is involved, one does not need to know the theory of relativity. In other words, in this standard approach, introducing as an only modification to the classical case of the relativistic mass, the time differs from space. We have no mixture of positions and time. This essential point has never received attention in the physical community.
Another difficulty is the confusion surrounding the combination of non-parallel velocities in special relativity. Let us consider the composition of two velocities when they are not lying on a straight line. In conventional particle tracking (similar to the Newton’s mechanics), one composes them geometrically by the method of a parallelogram with velocities regarded as the usual vectors. In relativistic kinematics, the composition of the velocities is studied analytically from the composition of Lorentz boosts. At the same time it is common knowledge that Lorentz boosts alone do not make a group. We must conclude that when we are dealing with transformations and a set of these transformations does not form a group we have no legal theory. Authors of textbooks believe that the relativity principle requires that Einstein’s velocity addition law gives the composition law of a group. The non-associativity of the composition law of the non-colinear velocities is not widely known. Many experts who learned the theory of relativity using textbooks will find this statement disturbing. The popular belief remains that Einstein’s law of addition of velocities is associative, although it is an exercise in elementary algebra to prove non-associativity  [^61] The non-commutativity and non-associativity of the relativistic composition of non-collinear velocities was first recognized by Mocanu [65] and latter resolved by Ungar [66], by showing role played by the Wigner rotation in the transformation.. It is therefore impossible to get covariant trajectory, $\vec{x}_{cov}(t)$, if we have nothing but Lorentz boosts. Something else has to be added. Lorentz boosts do not form a group due to the presence of the Wigner rotation. It comes out quite natural that the breakdown of commutativity and associativity in Einstein’s addition turns out to be Wigner rotation.
@@ 16.7 Convention-Invariant Particle Tracking
So far we have considered the motion of a particle in three-dimensional space using the vector-valued function $\vec{x}(t)$. We have a prescribed curve (path) along which the particle moves. The motion along the path is described by $l(t)$, where $l$ is a certain parameter (in our case of interest the length of the arc). Note the difference between the notions of path and trajectory. The trajectory of a particle conveys more information about its motion because every position is described additionally by the corresponding time instant. The path is rather a purely geometrical notion. Complete paths or their parts may consist of, e.g., line segments, arcs, circles, and helical curves. If we take the origin of the (Cartesian) coordinate system and connect the point to the point laying on the path, then the creating vector will be a position vector $\vec{x}(l)$. The derivative of a vector is the vector tangent to the curve described by the radius vector $\vec{x}(l)$. The sense of the $d\vec{x}(l)/dl$ is determined by the sense of the curve arc $l$.
We already know from our discussion in Chapter 2 that the path $\vec{x}(l)$ has exact objective meaning i.e. it is convention-invariant. The components of the momentum four-vector $mu=(\mathcal{E}/c,\vec{p})$ have also exact objective meaning. In contrast to this, and consistently with the conventionality intrinsic in the velocity, the trajectory $\vec{x}(t)$ of the particle in the lab frame is convention-dependent and has no exact objective meaning.
@@
These three equations correspond exactly to the equations for the components of the position vector that can be found using the non-covariant particle tracking approach, and $\vec{x}(l)_{cov}$ is exactly equal to $\vec{x}(l)$ as it must be. The point is that both approaches describe correctly the same physical reality and since the curvature radius of the path in the magnetic field, and consequently, the three-momentum, has an objective meaning (i.e. is convention-invariant), both approaches yield the same physical results.
17 Relativistic Spin Dynamics
We would like to show that while learning relativistic particle dynamics we simultaneously learned about a large number of other subjects. Let us consider the example of the motions of spin in a given field. Spin dynamics equations can be expressed as tensor equations in Minkowski space-time. We shall limit ourselves to the case of a particle with a magnetic moment $\vec{\mu}$ in a microscopically homogeneous electromagnetic field. Evidently, the torque affects only the spin and the force affects only the momentum. It follows that the motion of the system as a whole in any frame is determined entirely by its charge, independent of magnetic dipole moment. This part of the motion has been treated in the previous chapter. We need to consider spin motion only now.
17.1 Magnetic Dipole at Rest in an Electromagnetic Field
Let us first consider the spin precession for a nonrelativistic charged particle. The proportionality of magnetic moment $\vec{\mu}$ and angular momentum $\vec{s}$ has been confirmed in many ”gyromagnetic” experiments on many different systems. The constant of proportionality is one of the parameters charactering a particular system. It is normally specified by giving the gyromagnetic ratio or $g$ factor, defined by $\vec{\mu}=ge\vec{s}/(2mc)$. This formula says that the magnetic moment is parallel to the angular momentum and can have any magnitude. For an electron $g$ is very nearly 2.
Suppose that a particle is at rest in an external magnetic field $\vec{B}_{R}$. The equation of motion for the angular momentum in its rest frame is $d\vec{s}/d\tau=\vec{\mu}\times\vec{B}_{R}=eg\vec{s}\times\vec{B}_{R}/(2mc)=\vec{\omega}_{s}\times\vec{s}$. In other words, the spin precesses around the direction of the magnetic field with the frequency $\omega_{s}=-eg\vec{B}_{R}/(2mc)$. In the same nonrelativistic limit the velocity processes around the direction of $\vec{B}_{R}$ with the frequency $\omega_{p}=-(e/mc)\vec{B}_{R}$: $d\vec{v}/d\tau=(e/mc)\vec{v}\times\vec{B}_{R}$. Thus, for $g=2$ spin and velocity precess with the same frequency, so that the angle between them is conserved.
17.2 Derivation of the Covariant (BMT) Equation of Motion of Spin
In seeking the equation for the spin motion, we shall be guided by the known dynamics in the rest frame and the known relativistic transformation laws. We emphasize that spin is defined in a particular frame (the rest frame). Therefore, to form expressions with known transformation behavior, we need to introduce a four-quantity related to the spin. A convenient choice is a four- (pseudo)-vector $S$ defined by the requirement that in the rest frame, its space-like components are the spin components, while the time-like component is zero. We shall call $S$ four-spin; when normalized by dividing by its invariant length, it will be called polarization four-vector. It is space-like, and therefore in no frame does it space-like part vanish.
The BMT equation [46] is the manifestly covariant equation of motion for a four-vector spin $S^{\alpha}$ in an electromagnetic field $F^{\alpha\beta}$:
where $u_{\mu}=dx_{\mu}/d\tau$ is the four-dimensional particle velocity vector. With Eq.(21), one has
The BMT equation is valid for any given inertial frame, and consistently describes, together with the covariant-force law, the motion of a charged particle with spin and magnetic moment. The covariant equation of spin motion for a relativistic particle under the action of the four-force $Q^{\mu}=eF^{\mu\nu}u_{\nu}$ in the Lorentz lab frame, Eq.(27), is a relativistic ”generalization” of the equation of motion for a particle angular momentum in its rest frame. Relativistic ”generalization” means that the three independent equations expressing the Larmor spin precession are embedded into the four-dimensional Minkowski space-time.
In Lorentz coordinates there is a kinematics constraint $S^{\mu}u_{\mu}=0$, which is the orthogonality condition of four-spin and four-velocity. Because of this constraint, the four-dimensional dynamics law, Eq.(27), actually includes only three independent equations of motion. Using the explicit expression for Lorentz force we find that the four equations Eq.(27) automatically imply the constraint $S^{\mu}u_{\mu}=0$ as it must be. To prove this we may point out that one has in every Lorentz frame $S^{0}=\vec{S}\cdot\vec{v}$. While $S^{0}$ vanishes in the rest frame, $dS^{0}/d\tau$ need not. In fact $d(S^{\mu}u_{\mu})/d\tau=0$ implies $dS^{0}/d\tau=\vec{S}\cdot d\vec{v}/d\tau$. The immediate generalization of $d\vec{s}/d\tau=eg\vec{s}\times\vec{B}_{R}/(2mc)$ and $dS^{0}/d\tau=\vec{S}\cdot d\vec{v}/d\tau$ to arbitrary Lorentz frames is Eq.(27) as can be checked by reducing to the rest frame. A methodological analogy with the relativistic generalization of Newton’s second law emerges.